OpenStudy (poetryprincess):

If c(x) = 4x – 2 and d(x) = x^2 + 5x, what is (c*d)(x)? a. 4x^3 + 18x^2 – 10x b. X^2 + 9x – 2 c. 16x^2 + 4x – 6 d. 4x^2 + 20x – 2

OpenStudy (poetryprincess):

Someone please help me.

OpenStudy (anonymous):

* means multiplication or composition?

OpenStudy (poetryprincess):

composition

OpenStudy (anonymous):

if it means multiplication then (c*d) (x) = c(x) * d(x) = (4x-2) (x^2+5x)

OpenStudy (anonymous):

okay, in that case

OpenStudy (anonymous):

(c*d) (x) = c [d(x)] = c[x^2+5x] now this statement simply means that when you had c(x) = 4x-2 then c(2) = 4(2) -2 OR c(7) = 4(7) -2 OR c(x+2) = 4(x+2) -2 whatever follows the C we just plug that part into the RHS of the given function So, C(x^2+5x) = 4(x^2+5x)-2 = 4x^2 +20 x -2

OpenStudy (anonymous):

is it clear?

OpenStudy (poetryprincess):

Ah yes I see. Thanks so much for explaining all of that. I really appreciate it!!

OpenStudy (anonymous):

You are Welcome :)